The tight binding energy dispersion (E-k) relation for electrons in a one-dimensional array of atoms having lattice constant 'a' and total length 'L' is :

E=E0−β−2γcoska

Where E0,β and γ are constants and k is the wave vector.
The effective mass of electrons in the band is given by?
1
Expert's answer
2021-09-27T09:02:52-0400

Given:


"E(k)=E_0-\beta-2\gamma\cos ka"

The effective mass of electron is given by

"m^{*}=\frac{\hbar^2}{d^2E(k)/dk^2}"

Hence

"m^{*}=\frac{\hbar^2}{a^2*2\gamma \cos ka}=\frac{\hbar^2}{a^2(E_0-E+\beta)}"